OpenStudy (anonymous):

A circular swimming pool has a diameter of 12m. The sides are 3m high and the depth of the water is 1.5m. How much work (in Joules) is required to pump all of the water over the side? (The acceleration due to gravity is 9.8 m/s^2 and the density of water is 1000 kg/m^3

OpenStudy (anonymous):

weight of water = m*g = pi*(12/2)^2*1.5 * 1000 * 9.8 kg-m/ s² mass= volume*density = pi*(12/2)^2*1.5 * 1000 kg total volume of water = pi*(12/2)^2*1.5 m force = weight displacement = 3 m work = force*displacement = weight*displacement work =pi*(12/2)^2*1.5 * 1000 * 9.8 *3 kg-m²/ s² (1 kg-m²/ s² =1 J)

OpenStudy (anonymous):

Hope this helps